OpenStudy (anonymous):

Christina Hercher borrowed $50,000 on a 90 day, eight percent note. Christina paid $3,000 toward the note on day 40. On day 60 she paid an additional $4,000. Using the U.S. Rule, Christina's adjusted balance after the first payment is

ganeshie8 (ganeshie8):

1) calculate the interest amount for 40 days 2) add the interest to principal 3) subtract 3000 from the "above effective principal" 1) interest for 40 days on $50,000 = \[ \frac{.08}{360}*40*50,000 = 444.44\] 2) effective principal with interest until 40th day = 50,000 + 444.44 = 50,444.44 3) adjusted balance after first payment of $3000 = 50,444.44 - 3000 = 47,444.44

OpenStudy (anonymous):

thanks for the answer, but, why is it over 360 not 365? thanks again.